Time period of a planet whose semi major axis is 3.5 × 10⁹ m is found to be 100 years. Find the time period of a planet whose semi major of axis is twice of it.

Physics>Gravitation
Answer #1

According to the third Kepler's law,

\displaystyle \frac{a_1^3}{T_1^2} = \frac{a_2^3}{T_2^2} \;\; \Rightarrow \; \Big(\frac{T_2}{T_1}\Big)^2 = \Big(\frac{a_2}{a_1}\Big)^3

where T is a period and a is semi-major axis.

It is given that a_2=2a_1, so

\displaystyle \Big(\frac{T_2}{T_1} \Big)^2 = \Big(\frac{2a_1}{a_1} \Big)^3 = 2^3 = 8

\displaystyle T_2^2 = 8 T_1^2

\displaystyle T_2 = \sqrt{8} \,T_1 = 2.828 \cdot 100 = 282.8 \; \textrm{years}


answered by: Hans